# CloudFix: Automated Policy Repair for Cloud Access Control Policies Using Large Language Models

### v1

@@ III-D Repair Synthesizer (LLM)
We leverage an LLM to generate candidate repairs directly from misclassified requests and fault localization findings. The Repair Synthesizer LLM produces a candidate policy, which is then passed to the Goal Validator for verification. If the Goal Validator determines that the policy correctly classified the request specification, the policy will be returned as a successful repair. However, due to the complexity of the requests and the policies themselves, some faulty policies are difficult to fully repair in a single iteration. The repair cycle continues until the Goal Validator determines that the candidate policy correctly classifies the given must-allow and must-deny request specifications, or the predefined iteration threshold is reached. The candidate policy which achieved the highest classification accuracy is returned as a partially repaired policy.
@@
1:Misclassified requests $R$, policy $P$ with statements $S$
@@
8:  if $r.\mathrm{expected}=\mathrm{Deny}$ and $r.\mathrm{actual}=\mathrm{Allow}$ then
9:   $A_{r}\leftarrow\{\,s\in S^{+}\mid\mathrm{SMTSolver}(r,\{s\})=\mathrm{SAT}\,\}$
10:   $L\leftarrow L\cup\{(r,A_{r})\}$ $\triangleright$ explicit allow
11:  else if $r.\mathrm{expected}=\mathrm{Allow}$ and $r.\mathrm{actual}=\mathrm{Deny}$ then
12:   $D_{r}\leftarrow\{\,s\in S^{-}\mid\mathrm{SMTSolver}(r,P^{\star}\cup\{s\})=\mathrm{UNSAT}\,\}$
13:   if $D_{r}\neq\emptyset$ then
14:     $L\leftarrow L\cup\{(r,D_{r})\}$ $\triangleright$ explicit deny
15:   else
16:     $L\leftarrow L\cup\{(r,\emptyset)\}$ $\triangleright$ implicit deny
17:   end if
@@ IV Dataset Collection and Request Generation
CloudFix works to repair cloud access control policies that are written in cloud access control languages. However there is currently no public dataset specifically curated for policy repair, which makes research in this area difficult. This is because policies model access control within organizations and if these policies were made public they would could create security vulnerabilities. To address this gap, we curate real-world AWS IAM policies from the AWS re:Post forum and make them available to the research community. An important component of policy repair is the availability of representative request sets, but obtaining such a dataset is non-trivial because AWS IAM request logs are also kept private. To overcome these gaps, we algorithmically generate synthetic requests directly from the structure of the policy by varying the proportions of must-allow and must-deny requests. Below, we discuss the details of dataset collection and request generation.
@@ IV-A Policy Collection
The policy collection process involves three steps: first scraping policies from public sources, then detecting and categorizing valid JSON policies, and finally normalizing and repairing formatting inconsistencies. To collect a sufficiently large dataset of IAM policies, we sourced user posts from AWS re:Post. In this community forum, AWS users share questions and answers regarding AWS services and products. We specifically targeted posts containing IAM policies that either failed to function as intended or required correction. In total we scraped over 9,968 IAM-related posts. We then performed policy detection and categorization. Each post body was scanned for JSON objects resembling IAM policies. Detection required identifying bracketed structures containing keys such as Effect, Action, or Statement. All collected policies were then tested for syntactic correctness using Quacky [eiers2022quacky], a quantitative policy analyzer. Quacky revealed that many of the policies exhibited structural errors or formatting inconsistencies. To correct these issues, we implemented a naive syntactic repair tool to ensure we had a corpus of policies which adhered to standard IAM policy conventions. In total, we collected 282 faulty policies (syntactically correct but needing to be repaired) and 45 repaired policies (policies that were community-accepted as repairing the original broken policy). We summarize the statistics of our dataset in Table I.
TABLE I: Summary statistics of the AWS IAM policy dataset.
@@
   Allow (Count / %) 400   (89.7%)
   Deny (Count / %) 44   (9.9%)
@@ V Experimental Design
The goal of our evalution is to answer the following research questions:
RQ1: Can CloudFix repair cloud access control policies?
RQ2: How effective is Fault Localization for policy repair?
RQ3: How does accuracy vary across LLMs with FL-guided Repair?
RQ4: How well do the repairs generated by CloudFix generalize to requests not seen during the repair process?
RQ5: Can CloudFix repair policies with requests generated from natural language descriptions of faulty policies?
@@ V-A Experiment Setup
All experiments were run using a Rocky Linux 9.3 computer cluster (Blue Onyx). Each job was allocated a single NVIDIA L40S GPU with 46 GB of memory, two Intel(R) Xeon(R) Platinum 8562Y+ CPU cores, and 251 GB of system RAM. All experiments were executed with a 64 GB memory limit and a single GPU per task.
@@ V-B Models
We evaluate our approach using four open-source instruction-tuned models of comparable size. Specifically, we used CodeLLaMA-7B-Instruct, a model trained for code synthesis and reasoning; Granite-3.3-8B-Instruct, an IBM model optimized for instruction following in software engineering tasks; DeepSeek-Coder-7B-Instruct, a model finetuned for code completion and reasoning, designed for software repair and program synthesis, and Llama3.2-3B-Instruct, an instruction-tuned text-only model that supports a maximum of 128K context window. We maintain similar prompts and model hyperparameters for all models used. All models were executed through Hugging Face’s transformers pipeline [wolf2019huggingface].
@@ V-C Repair Approaches
We employ two repair approaches in our experiments. The first is a Baseline repair approach which uses a basic prompt to instruct the LLM to repair the policy. The Baseline receives the complete faulty policy, the requests that must be allowed and denied, and a generic repair instruction, where the model needs to identify faulty statements within the policy and repair them. The second is the Fault Localization guided approach, which is the crux of our framework and described in detail in Section III. We maintain the same user and system prompts for both approaches to ensure that the quality of the prompts does not confound the results.
Another important factor that influences repair accuracy is the size of the request. We use Algorithm 4 to generate 10, 20, 30, and 50 request sizes. To increase the complexity of the requests, we ensure that each request targets a unique resource–resource-action-condition-principal pair. Furthermore, our dataset exhibits a higher proportion of must-allow requests compared to must-deny requests, as shown in Table I. This distribution reflects realistic cloud usage patterns where most access requests are operations that must be allowed by access control policies.
@@ V-D Evaluation Metric
Accuracy. We evaluate our approach using accuracy, which is defined as the percentage of request specifications that are correctly classified by the policy. $\text{Accuracy}=\frac{R_{c}}{R_{\text{total}}}\times 100\%$, where $R_{c}$ is the number of correctly classified requests and $R_{\text{total}}$ is the total number of requests. We define three accuracy categorys: 100% (complete repair), 80-99% (moderate repair), and <80% (failed repair).
@@ V-E Experiments
RQ1: Can CloudFix repair cloud access control policies? Our first research question aims at testing whether our framework can automatically repair faulty policies such that all requests in the given specification are correctly classified.
Setup. To answer this question, we evaluate our framework on our dataset of 282 AWS IAM policies. For each policy, we generate must-allow and must-deny requests of 10, 20, 30, and 50 using Algorithm 4.
Results. Table III show that FL achieves 84.0% full repair rate on 10 requests compared to 48.2% for Baseline, and maintains 54.3% complete repair on 30 requests while Baseline drops to 22.3%. We see a performance degradation for 50 request sizes in both Baseline and Fault Localization, where the Baseline repairs 0 policies completely, while Fault Localization generates 7 completely repaired policies. The results demonstrate that FL consistently outperforms the Baseline as the request size increases, and gracefully degrades as Baseline performance dramatically decreases for higher request sizes.
We further performed a statistical significance analysis on the obtained results using a two-tailed t-test to compare the mean repair accuracies between the two approaches. The p-values below 0.001 for request sizes 10, 20, and 30 show strong statistical evidence (>99.9% confidence, see Table II, green cells) that FL genuinely outperforms Baseline. However, the p-value of 0.056 for request size 50 fails to meet the significance threshold of 0.05; meaning, the 3.85% improvement could be due to random chance rather than Fault Localization.
RQ2: How effective is Fault Localization for policy repair? Next, our second research question aims to determine whether FL improves repair accuracy. We compare our two approaches: (1) Baseline, which uses simple prompt consisting of only the faulty policy and requests, and (2) FL-guided repair, which additionally provides the fault localization information discussed in Section III.
@@
Results. Table III shows the distribution of repair accuracy across different request sizes. The FL-guided repair demonstrates significant improvements over the Baseline for request sizes 10, 20, 30, and 50. For request size 10, FL generates 237 complete repairs compared to only 136 for the Baseline. This advantage continues with FL achieving 170 complete repairs for 20 requests, 153 complete repairs for 30 requests, and 7 complete repairs for 50 requests.
For request size 10, the Baseline generates 101 policies in the <<80% category, while FL reduces this to only 19 policies. Similarly, for 30 requests, the Baseline generates 100 policies with <<80% accuracy compared to just 30 with FL. However, the Baseline achieves no 100% repairs but only 67 policies in the 80–99% range, while FL only achieves 7 complete repairs and produces 175 policies with <<80% accuracy compared to the 217 of the Baseline. We observed that as the request size increases, the repair task may exceed the LLM’s capability to integrate FL information effectively. We believe that the accuracy degradation at larger request sizes indicates that, while FL is highly effective for moderately complex repair tasks, additional techniques may be needed to maintain its benefits when the number of request sizes grows substantially.
TABLE II: Statistical Significance: Repair Accuracy of Baseline and FL
@@
p << 0.001; not significant: p >> 0.05.
Figure 4: Average repair accuracy of the Baseline and Fault Localization (FL) approaches against the initial faulty policies across varying request set sizes.
@@
Baseline Fault Localization Baseline Fault Localization Baseline Fault Localization Baseline Fault Localization
100% Fix 136 (48.22%) 237 (84.04%) 149 (52.83%) 170 (60.30%) 63 (22.34%) 153 (54.30%) 0 (0.00%) 7 (2.50%)
80–99% 45 (16.00%) 26 (9.22%) 62 (22.00%) 69 (24.50%) 119 (42.20%) 99 (35.10%) 67 (23.80%) 100 (35.46%)
<<80% 101 (35.82%) 19 (6.74%) 71 (25.20%) 43 (15.25%) 100 (35.50%) 30 (10.64%) 215 (76.24%) 175 (62.10%)
TABLE IV: Repair Time for Baseline and FL Across Request Sizes
@@
Baseline Fault Localization Baseline Fault Localization Baseline Fault Localization Baseline Fault Localization
Avg Iteration. 3.14 1.80 2.93 2.85 4.13 3.08 5.00 4.92
Avg Total Time Per Policy (s) 370.3 136.5 667.2 456.3 827.1 547.3 1778.6 1306.9
Avg SMT Time Per Policy (s) 2.98 0.27 8.58 0.44 11.05 1.11 8.26 3.92
Avg LLM Time Per Policy (s) 367.36 136.20 658.66 455.86 816.02 546.20 1770.29 1302.98
@@
RQ3: How does accuracy vary across LLMs with FL-guided repair? Following our evaluation of repair accuracy across request sizes, we now explore the performance of LLMs in repairing access control policies.
@@
Setup. We employed the instruction-tuned variants of CodeLlama-7B-Instruct, Llama-3.2-3B-Instruct, DeepSeek-Coder-7B-Instruct-v1.5, and Granite-3.3-8B-Instruct. We prompted all models using the Fault Localization-guided prompting strategy with a request size of 30 and for all of the 282 access control policies.
@@
Results. Figure 5 shows the overall repair accuracy of each model. Notably, CodeLlama achieves the highest overall accuracy of 92.3%, followed by Llama of 63.7%. DeepSeek-Coder and Granite had a relatively similar overall repair accuracy of 54.0% and 53.7%. Table V further shows the accuracy breakdown. CodeLlama was able to repair 54.3% of the 282 policies. The low accuracy in the rest of the models is best explained in the third category, which shows an increasing number of failed repairs, going from left to right.
@@
TABLE V: FL Repair Accuracy Comparison Across Different LLMs (Request Size 30)
Category Prompt CodeLlama Llama DeepSeek-Coder Granite
100% Fix FL-guided 153 (54.3%) 19 (6.7%) 37 (13.1%) 39 (13.8%)
80–99% FL-guided 99 (35.1%) 55 (19.5%) 32 (11.3%) 29 (10.3%)
<<80% FL-guided 30 (10.6%) 208 (73.8%) 213 (75.5%) 214 (75.9%)
@@
RQ4: How well do the repairs generated by CloudFix generalize to requests not seen during the repair process? The goal of this experiment is to test whether our repair framework can identify and generalize to semantically similar access patterns rather than overfitting to specific requests. A motivating example for this experiment involves multiple requests targeting similar resources,
(s3:getObject,arn:aws:s3:::category/data-1)
(s3:getObject,arn:aws:s3:::category/data-2)
(s3:getObject,arn:aws:s3:::category/data-3)
which could be grouped together to allows access to
(s3:getObject,arn:aws:s3:::category/data-*)
Enumerating similar requests this way has undesired effect. It challenges the LLM’s ability to generalize. As a result, simpler models might try to patch the policy by adding a specific rule for each file. While this works to correctly classify the provided requests, it fails for semantically related other requests and requires the models to reason more. Therefore, we test whether LLMs can repair policies in a way that generalizes to semantically similar access requests.
Setup. To answer RQ4, we evaluate the generalizability of CloudFix. To enumerate the requests, we first pick one misclassified request from the request sets of each policy and enumerate its resource into 10 new variants, similar to those conforming to the pattern shown above. This generated a total of 20 requests. Next, for each policy, we generated the repair using both the Baseline and FL-guided approaches on these 20 requests. To evaluate the repair on unseen requests, we then synthesized 15 variant requests from the same requests and evaluated whether the repaired policies correctly classified these unseen but semantically similar requests.
@@
Results. Figure 6 presents the repair and generalization accuracy for both approaches. The left panel shows that FL-guided repair achieves 82.8% (172 complete repairs) accuracy compared to the Baseline’s 79.9% (147 complete repairs), showing a 2.9% improvement over the initial 20.9% accuracy of the faulty policies. We can see from the right panel of the figure that the Baseline achieves 79.9% accuracy on unseen request variants, and FL achieves 84.9%. Interestingly, the Baseline generalization accuracy matches its repair accuracy (79.9%). This suggests that when the Baseline generates complete repairs, it produces repairs that generalize well to semantically similar requests. The FL approach shows a stronger generalization, achieving 84.9% on variants compared to its 82.8% repair accuracy.
@@
Figure 5: An overall repair accuracy of four different LLMs using the FL-guided prompting strategy with a request size of 30.
@@
Figure 6: The left panel compares the repair accuracy of the Baseline (79.9%) and FL-guided (82.8%) methods against the initial faulty policy (20.9%). The right panel evaluates how these repairs generalize to unseen requests, showing the FL-guided approach (84.9%) can generate more generalizable repairs than the Baseline (79.9%).
RQ5: Can CloudFix repair policies with requests generated from natural language descriptions of faulty policies Finally, we test whether CloudFix can completely repair policies using requests generated by LLMs.
@@
Setup. To answer RQ5, we select 10 faulty policies from our dataset. As mentioned in Section IV, our AWS IAM dataset contains 45 policies that have faulty policy: the original, incorrect policy posted by a user experiencing access control issues, ground truth policy: the corrected policy provided by an AWS Cloud Engineer in the forum, and user intent: the natural language description of the request specifications or problem as described by the user. We first use GPT-5, accessed via it’s web interface, to synthesize requests. We follow the following procedure: for each policy, we prompted the model with the user intent, faulty policy, and ground truth policy and instruct it to generate 5 up to 10 valid requests. The criteria to generate a valid request were that they must be misclassified by the faulty policy, correctly classified by the ground truth, and encapsulate the user’s intent. We then used Algorithm 2 to validate that the requests were misclassified by the faulty policy and correctly classified by the ground truth policy. To validate if the requests align with the user’s expressed intent, we manually verified to ensure correctness. Next, we used CodeLlama-7B-Instruct to repair the 10 faulty policies with a maximum of 5 iterations and a target accuracy of 100%.
@@
Results. Figure 7 shows that the Baseline approach achieves 85.2% repair accuracy, while FL achieves 86.4%, both showing improvements over the 46.7% initial accuracy. This demonstrates that our framework can be used to generate policy repairs using requests generated by LLMs from natural langauge descriptions of the users intents.
@@
Figure 7: The results for CloudFix on requests generated from natural language descriptions. The marginal difference between the two methods suggests that the requests generated by the LLM from user intent were effective at encapsulating the nuanced semantic difference.

### v2

@@ III-D Repair Synthesizer (LLM)
We leverage an LLM to generate candidate repairs directly from misclassified requests and Fault Localization findings. The Repair Synthesizer LLM produces a candidate policy, which is then passed to the Goal Validator for verification. If the Goal Validator determines that the policy correctly classified the request specification, the policy will be returned as a successful repair. However, due to the complexity of the requests and the policies themselves, some faulty policies are difficult to fully repair in a single iteration. The repair cycle continues until the Goal Validator determines that the candidate policy correctly classifies the given must-allow and must-deny request specifications, or the predefined iteration threshold is reached. The candidate policy, which achieved the highest classification accuracy, is returned as a partially repaired policy.
@@
1:Request set $R$, policy $P$ with statements $S$
@@
8:  if $\textsc{ValidateGoal}(P,\{r\})=\mathrm{FAIL}$ then
9:   if $r.\mathrm{expected}=\mathrm{Deny}$ and $r.\mathrm{actual}=\mathrm{Allow}$ then
10:     $A_{r}\leftarrow\{\,s\in S^{+}\mid\mathrm{SMTSolver}(r,\{s\})=\mathrm{SAT}\,\}$
11:     $L\leftarrow L\cup\{(r,A_{r})\}$ $\triangleright$ Case 1: explicit allow
12:   else if $r.\mathrm{expected}=\mathrm{Allow}$ and $r.\mathrm{actual}=\mathrm{Deny}$ then
13:     $D_{r}\leftarrow\{\,s\in S^{-}\mid\mathrm{SMTSolver}(r,P^{\star}\cup\{s\})=\mathrm{UNSAT}\,\}$
14:     if $D_{r}\neq\emptyset$ then
15:      $L\leftarrow L\cup\{(r,D_{r})\}$ $\triangleright$ Case 2: explicit deny
16:     else
17:      $L\leftarrow L\cup\{(r,\emptyset)\}$ $\triangleright$ Case 3: implicit deny
@@ IV Dataset Collection and Request Generation
CloudFix works to repair cloud access control policies that are written in cloud access control languages. However, there is currently no public dataset specifically curated for policy repair, which makes research in this area difficult. This is because policies model access control within organizations, and if these policies were made public, they could create security vulnerabilities. To address this gap, we curate real-world AWS IAM policies from the AWS re:Post forum and make them available to the research community. An important component of policy repair is the availability of representative request sets, but obtaining such a dataset is non-trivial because AWS IAM request logs are also kept private. To overcome these gaps, we algorithmically generate synthetic requests directly from the structure of the policy by varying the proportions of must-allow and must-deny requests.
@@ IV-A Policy Collection
The policy collection process involves three steps: first, scraping policies from public sources, then detecting and categorizing valid JSON policies, and finally normalizing and repairing formatting inconsistencies. To collect a sufficiently large dataset of IAM policies, we sourced user posts from AWS re:Post. In this community forum, AWS users share questions and answers regarding AWS services and products. We specifically targeted posts containing IAM policies that either failed to function as intended or required correction. We scraped 9,968 IAM-related posts and scanned each for JSON structures containing IAM policy keys such as Effect, Action, or Statement. We validated all collected policies using Quacky[8], a quantitative policy analyzer, which revealed widespread structural errors and formatting inconsistencies. After applying a naive syntactic repair tool to enforce standard IAM conventions, we obtained 282 faulty policies (syntactically valid but semantically broken) and 45 corresponding community-accepted repairs. Table I summarizes our dataset.
TABLE I: Summary Statistics of the AWS IAM Policy Dataset
@@
     Allow 400 (89.7%)
     Deny 44 (9.9%)
Percentages are relative to total statements.
@@ V Experimental Design
We address the following research questions:
RQ1: How effective is Fault Localization at improving LLM-based policy repair compared to a baseline?
RQ2: How does request size affect repair accuracy?
RQ3: How does repair accuracy vary across LLMs?
RQ4: Can CloudFix generate repairs that generalize to semantically similar requests beyond those provided during the repair process?
RQ5: Can CloudFix repair policies when request specifications are synthesized from natural language descriptions of user intent?
@@ V-A Experiment Setup
All experiments were run using a Rocky Linux 9.3 computer cluster (Blue Onyx). Each job was allocated a single NVIDIA L40S GPU with 46 GB of memory, two Intel(R) Xeon(R) Platinum 8562Y+ CPU cores, and 256 GB of system RAM. All experiments were executed with a 64 GB memory limit and a single GPU per task.
@@ V-B Models
We evaluate our approach using four open-source instruction-tuned models of comparable size. Specifically, we used CodeLLaMA-7B-Instruct, a model trained for code synthesis and reasoning; Granite-3.3-8B-Instruct, an IBM model optimized for instruction following in software engineering tasks; DeepSeek-Coder-7B-Instruct, a model finetuned for code completion and reasoning, designed for software repair and program synthesis, and Llama3.2-3B-Instruct, an instruction-tuned text-only model that supports a maximum of 128K context window. We maintain similar prompts and model hyperparameters for all models used. All models were loaded and run via Hugging Face’s transformers pipeline [27].
@@ V-C Repair Approaches
We employ two approaches: first, a baseline repair approach that uses the basic prompt to instruct the LLM to repair a faulty policy. The baseline receives the complete faulty policy, the requests that must be allowed and denied, and a generic repair instruction, where the model needs to identify faulty statements within the policy and repair them. Second, a Fault Localization guided approach, described in detail in Section III, which represents the core contribution of CloudFix. The baseline serves as an ablation to isolate the impact of Fault Localization. We maintain the same user and system prompts for both approaches to ensure that the quality of the prompts does not confound the results. Request size is another key factor that influences repair accuracy. Using Alg. 4, we generate request sets of size 10, 20, 30, and 50, where each request targets a unique resource–action–condition–principal combination to increase complexity.
@@ V-D Evaluation Metric
Accuracy. We evaluate our approach using accuracy, which is defined as the percentage of request specifications that are correctly classified by the policy. $\text{Accuracy}=\frac{R_{c}}{R_{\text{total}}}\times 100\%$, where $R_{c}$ is the number of correctly classified requests and $R_{\text{total}}$ is the total number of requests. We define three accuracy categories: $100\%$ (complete repair), $80-99\%$ (moderate repair), and $80\%$ (failed repair).
@@ V-E Experiments
RQ1: How effective is Fault Localization at improving LLM-based policy repair compared to a baseline? Our first research question evaluates whether Fault Localization guidance improves the ability of LLMs to repair faulty access control policies compared to a baseline approach.
Setup. To answer RQ1, we evaluate our framework on 282 AWS IAM policies, generating 10, 20,30, and 50 must-allow and must-deny requests per policy using Alg. 4.
Results. Table III shows that Fault Localization achieves 84.0% full repair rate on 10 requests compared to 48.2% for baseline, and maintains 54.3% complete repair on 30 requests while baseline drops to 22.3%. We see a performance degradation for 50 request sizes in both baseline and Fault Localization, where the baseline repairs 0 policies completely, while Fault Localization generates 7 completely repaired policies. The results demonstrate that Fault Localization consistently outperforms the baseline as the request size increases, and gracefully degrades as baseline performance dramatically decreases for higher request sizes.
We further performed a statistical significance analysis on the obtained results using a two-tailed t-test to compare the mean repair accuracies between the two approaches. The p-values below 0.001 for request sizes 10, 20, and 30 show strong statistical evidence (>99.9% confidence, see Table II, shaded rows) that Fault Localization outperforms the baseline. However, the p-value of 0.056 for request size 50 fails to meet the significance threshold of 0.05, meaning the 3.85% improvement could be due to random chance rather than the true impact of Fault Localization.
RQ2: How does request size affect repair accuracy?
Our second research question examines how repair accuracy changes as the number of requests increases. We compare our two approaches: (1) baseline, which uses a simple prompt consisting of only the faulty policy and requests, and (2) Fault Localization-guided repair, which additionally provides the Fault Localization information discussed in Section III.
@@
Results. Table III shows the distribution of repair accuracy across different request sizes. The Fault Localization guided repair demonstrates significant improvements over the baseline for request sizes 10, 20, 30, and 50. For request size 10, Fault Localization generates 237 complete repairs compared to only 136 for the baseline. This advantage continues with Fault Localization, achieving 170 complete repairs for 20 requests, 153 complete repairs for 30 requests, and 7 complete repairs for 50 requests.
For request size 10, the baseline generates 101 policies in the <<80% category, while Fault Localization reduces this to only 19 policies. Similarly, for 30 requests, the baseline generates 100 policies with <<80% accuracy compared to just 30 with Fault Localization. However, the baseline achieves no 100% repairs but only 67 policies in the 80–99% range, while Fault Localization only achieves 7 complete repairs and produces 175 policies with <<80% accuracy compared to the 215 of the baseline. We observed that as the request size increases, the repair task may exceed the LLM’s capability to integrate Fault Localization information effectively. The accuracy degradation at larger request sizes indicates that, while Fault Localization is effective for moderately complex repair tasks, additional techniques may be needed to maintain its benefits when the number of request sizes grows substantially.
TABLE II: Statistical Significance: Repair Accuracy of Baseline and Fault Localization
@@
***$p<0.001$. Shaded rows indicate statistical significance.
Figure 4: Average repair accuracy of Baseline and FL approaches across varying request sizes.
@@
Baseline FL Baseline FL Baseline FL Baseline FL
100% Fix 136 (48.2%) 237 (84.0%) 149 (52.8%) 170 (60.3%) 63 (22.3%) 153 (54.3%) 0 (0.0%) 7 (2.5%)
80–99% 45 (16.0%) 26 (9.2%) 62 (22.0%) 69 (24.5%) 119 (42.2%) 99 (35.1%) 67 (23.8%) 100 (35.5%)
<<80% 101 (35.8%) 19 (6.7%) 71 (25.2%) 43 (15.3%) 100 (35.5%) 30 (10.6%) 215 (76.2%) 175 (62.1%)
FL denotes Fault Localization; bold indicates the best complete repair rate for each request size.
TABLE IV: Repair Time for Baseline and Fault Localization Across Request Sizes
@@
Baseline FL Baseline FL Baseline FL Baseline FL
Avg. Iterations 3.14 1.80 2.93 2.85 4.13 3.08 5.00 4.92
Avg. Total Time (s) 370.3 136.5 667.2 456.3 827.1 547.3 1778.6 1306.9
Avg. SMT Time (s) 2.98 0.27 8.58 0.44 11.05 1.11 8.26 3.92
Avg. LLM Time (s) 367.36 136.20 658.66 455.86 816.02 546.20 1770.29 1302.98
@@
FL denotes Fault Localization; bold indicates faster convergence (fewer iterations).
RQ3: How does repair accuracy vary across different open-source LLMs? We now compare LLM performance on policy repair.
@@
Setup. We employed the instruction-tuned variants of CodeLlama-7B-Instruct, Llama-3.2-3B-Instruct, DeepSeek-Coder-7B-Instruct-v1.5, and Granite-3.3-8B-Instruct. We prompted all models using the Fault Localization guided prompt with a request size of 30 for all 282 access control policies.
@@
Results. Figure 5 shows the overall repair accuracy of each model. Notably, CodeLlama achieves the highest overall accuracy of 92.3%, followed by Llama of 63.7%. DeepSeek-Coder and Granite had a relatively similar overall repair accuracy of 54.0% and 53.7%. Table V further shows the accuracy breakdown. CodeLlama was able to repair 54.3% (complete repair rate) of the 282 policies . The low accuracy in the rest of the models is best explained in the third category, which shows an increasing number of failed repairs, going from left to right.
@@
TABLE V: Fault Localization Repair Accuracy Comparison Across Different LLMs (Request Size 30)
Category CodeLlama Llama DeepSeek-Coder Granite
100% Fix 153 (54.3%) 19 (6.7%) 37 (13.1%) 39 (13.8%)
80–99% 99 (35.1%) 55 (19.5%) 32 (11.3%) 29 (10.3%)
<<80% 30 (10.6%) 208 (73.8%) 213 (75.5%) 214 (75.9%)
Bold, shaded cells denote the best-performing model (CodeLlama); results use Fault Localization prompting.
@@
RQ4: Can CloudFix generate repairs that generalize to semantically similar requests beyond those provided during the repair process? This question tests whether repaired policies capture general access patterns rather than overfitting to specific requests. For example, requests targeting (s3:getObject,arn:aws:s3:::category/data1)
(s3:getObject,arn:aws:s3:::category/data2)
(s3:getObject,arn:aws:s3:::category/data3)
repeating up to dataN can be generalized to:
(s3:getObject,arn:aws:s3:::category/data*)
Without generalization, the LLM may add a specific rule for each request, correctly classifying provided requests but failing on similar unseen ones.
Setup. For each policy, we selected one misclassified request and generated 10 resource variants (e.g., data1 through data10), producing 20 total requests per policy. We repaired each policy using both approaches, then tested generalizability by evaluating the repaired policies on 15 held-out variant requests not provided during the repair process.
@@
Results. Figure 6 presents the repair and generalization accuracy for both approaches. The left panel shows that Fault Localization guided repair achieves 82.8% (172 complete repairs) accuracy compared to the baseline’s 79.9% (147 complete repairs), showing a 2.9% improvement over the baseline. The right panel shows that the baseline achieves 79.9% accuracy on unseen request variants, and Fault Localization achieves 84.9%. Interestingly, the baseline generalization accuracy matches its repair accuracy (79.9%). This suggests that when the baseline generates complete repairs, it produces repairs that generalize well to semantically similar requests. The Fault Localization approach shows stronger generalization, achieving 84.9% on variants compared to its 82.8% repair accuracy.
@@
Figure 5: Overall repair accuracy of four open-source LLMs using Fault Localization guided with a request size of 30.
@@
Figure 6: Left panel: compares the repair accuracy of the baseline (79.9%) and Fault Localization guided (82.8%) methods against the initial faulty policy (20.9%). Right panel: evaluates how these repairs generalize to requests not provided during the repair process, showing Fault Localization (84.9%) can generate more generalizable repairs than the baseline (79.9%).
RQ5: Can CloudFix repair policies when requests are synthesized from natural language descriptions of user intent? Our final research question examines whether CloudFix can repair policies when an LLM from natural language descriptions of user intent generates requests. In practice, administrators often know what access they want to allow or deny but may lack request specifications to test their policies. Manually crafting such specifications is tedious and prone to errors. If LLMs can translate natural language intent into concrete requests, CloudFix becomes applicable even when requests are unavailable.
@@
Setup. We select 10 faulty policies from our dataset. As mentioned in Section IV, our AWS IAM dataset contains 45 policies that have faulty policy (the original, incorrect policy posted by a user experiencing access control issues), ground truth policy (the corrected policy provided by an AWS Cloud Engineer), and user intent (the natural language description of the request specifications or problem as described by the user). We first use GPT-5, accessed via its web interface, to synthesize requests. We follow this procedure: for each policy, we prompted the model with the user intent, faulty policy, and ground truth policy and instructed it to generate 5 to 10 valid requests. The criteria to generate a valid request were that they must be misclassified by the faulty policy, correctly classified by the ground truth, and encapsulate the user’s intent. We then used Alg. 2 to validate that the requests were misclassified by the faulty policy and correctly classified by the ground truth policy. To validate if the requests align with the user’s expressed intent, we manually verified to ensure alignment. Next, we used CodeLlama-7B-Instruct to repair the 10 faulty policies with a max of 5 iterations and a 100% target accuracy.
@@
Results. Figure 7 shows that the baseline approach achieves 85.2% repair accuracy, while Fault Localization achieves 86.4%, both showing improvements over the 46.7% initial accuracy. This demonstrates that our framework can be used to generate policy repairs using requests generated by LLMs from natural language descriptions of the users’ intents.
@@
Figure 7: The results of CloudFix on requests generated from natural language descriptions. The marginal difference between the two methods suggests that LLM-generated requests from user intent effectively capture nuanced semantic differences.